### v1

# A Zeroth-Order Extra-Gradient Method For Black-Box Constrained Optimization

Abstract
Non-analytical objectives and constraints often arise in control systems, particularly in problems with complex dynamics, which are challenging yet lack efficient solution methods. In this work, we consider general constrained optimization problems involving black-box objectives and constraints. To solve it, we reformulate it as a min-max problem and propose a zeroth-order extra gradient (ZOEG) algorithm that combines the extra gradient method with a feedback-based stochastic zeroth-order gradient estimator. Then, we apply another coordinate gradient estimator to design the zeroth-order coordinate extra gradient algorithm (ZOCEG) to further improve efficiency. The theoretical analysis shows that ZOEG can achieve the best-known oracle complexity of $\mathcal{O}(d\epsilon^{-2})$ to get an $\epsilon$-optimal solution ($d$ is the dimension of decision space), and ZOCEG can improve it to $\mathcal{O}(d\epsilon^{-1})$. Furthermore, we develop a variant of ZOCEG, which applies block coordinate updates to enhance the efficiency of single-step gradient estimation. Finally, numerical experiments on a load tracking problem validate our theoretical results and the effectiveness of the proposed algorithms.
1   Introduction
In this paper, we consider a general constrained optimization problem in complex systems where the objective and constraint functions are deterministic but black-box, i.e., we only have access to their zeroth-order feedback (function values) rather than any higher-order information, such as gradients and Hessians. Specifically, we consider the following constrained optimization problem:
$\min_{\mathbf{x}\in\mathcal{X}}\ \phi_{0}(\mathbf{x})\ \ \text{s.t.}\ \phi_{j}(\mathbf{x})\leq 0,\ \forall j=1,2,...,d_{y},$ (1)
where $\mathcal{X}\in\mathbb{R}^{d_{x}}$ is a convex and compact set, and $\phi_{j}:\mathbb{R}^{d_{x}}\to\mathbb{R},\ \forall j=0,1,...,d_{y}$ are convex functions. They have no analytical forms, but can be evaluated at any query point $x\in\mathbb{R}^{d_{x}}$. Such optimization problems with black-box dynamics arise in diverse domains, including optimal control [13], simulation optimization [11], machine learning [30], power systems optimization [5], etc. In these problems, the objective and constraints usually lack analytical models due to the high complexities of the dynamics, expensive computation, and privacy issues [17]. Therefore, we can only access their zeroth-order feedback via measurements or simulations. For instance, we can hardly model the dynamics of a distribution network in power systems when information about the topology and parameters is unavailable. As a result, we can only observe the feedback, such as the voltages and power flow, by performing the given input.
Generally, traditional gradient-based algorithms, which rely on gradient computation, cannot be directly applied to problems with black-box dynamics. This necessitates the use of derivative-free optimization methods. Among them, direct search approaches are widely adopted [15], including the simplex methods and directional direct search (DDS) methods. By constructing and improving a simplex according to the evaluated function values at the vertices, the Nelder-Mead simplex algorithm is one of the most popular derivative-free methods [26]. The DDS methods typically consist of a search step and a poll step. In the search step, a candidate set of points is evaluated to check the improvement in function values. In the poll step, neighboring points are explored when no improvement is detected in the search step. Various algorithms fall under this framework, primarily distinguished by how the poll directions are generated. Specifically, coordinate search uses elementary basis vectors for poll directions [9], while generalized pattern search (GPS) samples poll directions from a positive spanning set defined on a mesh [36], which allows more flexible and adaptive search directions [18]. Furthermore, the mesh adaptive direct search (MADS) method generalizes GPS to achieve improved performance by adapting the mesh dynamically [1]. Nevertheless, these methods are restricted to low-dimensional problems. Another line of research focuses on model-based approaches. These studies develop surrogate models to approximate the objective and constraint functions. The commonly used surrogates include polynomial models [6], Gaussian process [37], Radial Basis Function (RBF) [31], etc. The algorithm alternates between the following two steps until termination: (i) evaluating the objective values at sampled points and updating the surrogate model; and (ii) sampling new points based on the surrogate model and past data, a process that often involves solving subproblems [32]. However, their performance is highly dependent on the choice of surrogate models.
Many techniques have been developed in derivative-free optimization to deal with black-box constraints, mainly including penalty, filter, and approximation methods. For penalty methods, a merit function combining the objective function and the constraint violation is minimized [8]. Filter methods are inspired by bi-objective optimization, where each update seeks to decrease either the objective function value or the constraint violation [34]. For model-based approaches, approximation models are built for both the objective and constraint functions to handle the black-box constraints [33]. However, most studies mentioned above lack convergence analysis, and the few that establish convergence results only provide asymptotic analysis.
Zeroth-order optimization (ZO), another representative method in derivative-free optimization, has recently gained increasing attention due to its ease of implementation and capability to deal with high-dimensional problems [38]. The core idea behind ZO is to use the finite difference of function values to construct a gradient estimator and embed it into gradient-based algorithms [2]. The gradient estimation includes single-point, $2$-point, and $2d$-point approaches, where $d$ is the dimension of the decision variable. Wherein, the $2d$-point coordinate gradient estimator (CooGE) is the most intuitive one that estimates the partial gradient along each dimension asynchronously, but it is inefficient to generate a full gradient for high-dimensional problems [39, 21]. The single-point and $2$-point gradient estimators synchronously add perturbations sampled from isotropic distributions, e.g., the Gaussian distribution and uniform distribution on a unit sphere (leading to GauGE and UniGE, respectively), to all dimensions to estimate the full gradients [42, 10, 29, 35]. However, $2$-point estimators are usually preferred because the variance of single-point estimators is unbounded, thus leading to poor performance in optimization [7]. By combining zeroth-order gradient estimators with gradient-based algorithms, e.g., gradient descent and mirror descent, ZO can generally achieve similar convergence rates as their first-order counterparts in unconstrained optimization, differing only by a factor of $d$ [19].
When it comes to constrained optimization, the majority of existing ZO research focuses on problems with black-box objectives but simple and explicit constraint sets. These white-box constraints allow efficient Euclidean or Bregman projections or admit linear-minimization oracles, which enable efficient approaches to maintain feasibility at each iteration. For instance, integrating zeroth-order gradient estimators with projected gradient descent yields the ZO-PSGD algorithm [12], which can be further extended to the mirror descent framework [7, 41]. In addition, projection-free methods have been developed based on the Frank-Wolfe method [4, 40]. To deal with black-box constraints (as in (1)), primal-dual frameworks are widely employed to solve the derived min-max problems [16]. By combining constraint extrapolation techniques with zeroth-order stochastic gradient estimators, the oracle complexity (number of function value queries) of $\mathcal{O}(d^{2}\epsilon^{-2})$ is achieved in [30], where $\epsilon$ is the target error. Besides, the zeroth-order gradient descent ascent (GDA) algorithm was developed to solve the min-max problems [20]. A variant of zeroth-order GDA, called zeroth-order optimistic GDA, is proposed, and the oracle complexity of $\mathcal{O}(d^{4}\epsilon^{-2})$ is established [22]. However, the algorithms tailored to deal with problem (1) are still scarce, and the existing ones suffer from slow convergence due to the large variance of stochastic gradient estimation.
To address the challenges of ZO with black-box constraints, we design a zeroth-order extra-gradient (ZOEG) algorithm that integrates the $2$-point UniGE and a variant of GDA, called extra-gradient (EG) [27], in this work to solve the problem (1). In addition, we perform finite-sample analysis of ZOEG to achieve improved dependence on the dimension $d$. Furthermore, to address the variances of the $2$-point uniform gradient estimator, we leverage the $2d$-point CooGE with the EG method and design the zeroth-order coordinate extra-gradient (ZOCEG) algorithm to achieve improved oracle complexity. We further extend ZOCEG to zeroth-order block coordinate extra-gradient (ZOBCEG) by randomly selecting subsets of coordinates for updates at each iteration, thereby enhancing the single-step computational efficiency. The main results are summarized in Table 1, and the contributions of this paper are summarized as follows.
  • We apply the primal-dual framework to solve problem (1) and reformulate it into a min-max problem. Then, we design the ZOEG algorithm by embedding the $2$-point UniGE into the EG method to solve the problem without requiring any higher-order information. The theoretical analysis of ZOEG provides the oracle complexity $O(d\epsilon^{-2})$ to find an $\epsilon$-optimal solution with the best-known dependence on $d$.
  • To further address the large variance of $2$-point estimators, we improve the proposed ZOEG by applying the coordinate gradient estimators and propose a new algorithm, ZOCEG, which estimates the full gradient information with the $2d$-point CooGE. The theoretical analysis establishes the oracle complexity $O(d\epsilon^{-1})$ of ZOCEG. This result represents, to the best of our knowledge, the state-of-the-art iteration complexity for solving problem (1). We also extend ZOCEG to ZOBCEG to enhance the efficiency of gradient estimation.
  • We validate the proposed algorithms and theoretical results on a practical load tracking task in power systems, and compare our algorithms with existing algorithms. The proposed algorithms show superior convergence and also perform well in nonconvex problems.
Table 1: Comparison of complexity bounds.
Algorithm
Gradient
Estimator
Oracle
Complexity
OGDA-RR[22] 2-point GauGE $\mathcal{O}(d^{4}\epsilon^{-2})$
SZO-ConEX[30] 2-point GauGE $\mathcal{O}(d^{2}\epsilon^{-2})$
ZOEG (Ours) 2-point UniGE $\mathcal{O}(d\epsilon^{-2})$
ZOCEG (Ours) $2d$-point CooGE $\mathcal{O}(d\epsilon^{-1})$
ZOBCEG (Ours) $2\tau$-point CooGE $\mathcal{O}(d\epsilon^{-2})$
Note: $\tau$ denotes the block size of randomly selected coordinates (see detailed definition in Section 4).
The rest of the paper is organized as follows. Section II introduces some preliminaries of the methodology in our work. In Section III, we present the ZOEG algorithm and establish its finite-sample convergence. The design and convergence results of the ZOCEG algorithm and its extension are presented in Section IV. We verify the performance of the proposed algorithms via numerical experiments in Section V. Finally, Section VI concludes this paper.
2   Preliminaries
[…]
2.1   Problem formulation and assumptions
To address the constrained optimization problem (1), we first apply the primal-dual framework and derive the corresponding Lagrangian function as follows:
$f(\mathbf{x},\mathbf{y})=\phi_{0}(\mathbf{x})+\sum_{j=1}^{d_{y}}y_{j}\phi_{j}(\mathbf{x}),$
which is a convex-concave function, i.e., $f(\mathbf{x},\mathbf{y})$ is convex in $\mathbf{x}$ and concave in $\mathbf{y}$. Then, given the problem (1), we can instead consider the following general convex-concave min-max problem:
$\min_{\mathbf{x}\in\mathcal{X}}\max_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y}),$ (2)
where $\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$ is a convex constraint set. For notational simplicity, we denote $\mathbf{z}=(\mathbf{x},\mathbf{y})$, and $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, $d=d_{x}+d_{y}$. Define the operator $F(\mathbf{z}):=[\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y});-\nabla_{\mathbf{y}}f(\mathbf{x},\mathbf{y})]$. To evaluate the solution quality of problem (2), we define the duality gap for any $\mathbf{z}=(\mathbf{x},\mathbf{y})\in\mathcal{Z}$:
$\Delta(\mathbf{z})=f(\mathbf{x},\mathbf{y}^{*})-f(\mathbf{x}^{*},\mathbf{y}),$
where $\mathbf{z}^{*}=(\mathbf{x}^{*},\mathbf{y}^{*})$ denotes the min-max saddle point of $f(\mathbf{x},\mathbf{y})$. The existence of $\mathbf{z}^{*}$ is assumed throughout this research. Accordingly, a feasible point $\mathbf{z}\in\mathcal{Z}$ is said to be an $\epsilon$-saddle point of $f$ if $\Delta(\mathbf{z})\leq\epsilon$. Clearly, $\mathbf{z}$ is the saddle point of $f$ if $\Delta(\mathbf{z})=0$, which provides an optimal solution to problem (1). Similarly, we say a point $\mathbf{x}\in\mathbb{R}^{d_{x}}$ is $\epsilon$-optimal for problem (1) if $[\bm{\phi}(\mathbf{x})]_{+}\leq\epsilon$ and $\bm{\phi}_{0}(\mathbf{x})-\bm{\phi}_{0}^{*}\leq\epsilon$, where $\bm{\phi}=(\phi_{1},\dots,\phi_{d_{y}})^{T}$ denotes the vector of constraint functions, and $\phi_{0}^{*}=\phi_{0}(\mathbf{x}^{*})$ denotes the optimal objective value.
Then, we formally provide some mild assumptions on the problem (2), which will be used throughout the analysis.
Assumption 2.1.
The function $f(\mathbf{x},\mathbf{y})$ is continuously differentiable over $\mathcal{X}\times\mathcal{Y}$. Furthermore, $f(\mathbf{x},\mathbf{y})$ is convex-concave, i.e., convex in $x$ and concave in $y$.
Assumption 2.2.
The function $f$ is $G$-Lipschitz and $L$-smooth, i.e., for any $\mathbf{z}_{1},\mathbf{z}_{2}\in\mathcal{Z}$,
$\|\nabla f(\mathbf{z}_{1})-\nabla f(\mathbf{z}_{2})\|\leq L\|\mathbf{z}_{1}-\mathbf{z}_{2}\|,$
$\|f(\mathbf{z}_{1})-f(\mathbf{z}_{2})\|\leq G\|\mathbf{z}_{1}-\mathbf{z}_{2}\|.$
Assumption 2.3.
The set $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ is compact and convex, thus, we have $\tilde{D}=\max_{\mathbf{z},\mathbf{z}^{\prime}\in\mathcal{Z}}\|\mathbf{z}-\mathbf{z}^{\prime}\|\leq+\infty$
Assumption 2.1 can be equivalently imposed on the primal problem, i.e., $\bm{\phi}_{j}(\mathbf{x}),\forall j\in\{0,1,\cdots,d_{y}\}$ is continuously differentiable and convex. Assumption 2.2 imposes the Lipschitz continuity on $f(\mathbf{x},\mathbf{y})$ and its gradients, which is common in smooth optimization [20, 17]. As for Assumption 2.3, the boundedness of $\mathcal{X}$ is natural because $x$ is generally bounded in practical problems. Although $\mathbf{y}\geq 0$ is the Lagrange multiplier and unbounded, the boundedness of $\mathbf{y}^{*}$ is justified in [25] under the Slater condition. Therefore, we can construct a bounded set $\mathcal{Y}$ containing $\mathbf{y}^{*}$ to replace $\{\mathbf{y}\in\mathbb{R}^{d_{y}}|\mathbf{y}\geq 0\}$.
2.2   Zeroth-order gradient estimators
Note that we can only obtain the function values of $\phi_{0}(\mathbf{x})$ and $\phi_{j}(\mathbf{x})$ for all $j\in[d_{y}]$, but lack access to higher-order information such as gradients and Hessians. This limitation renders traditional first-order min-max optimization methods inapplicable. To address this, we turn to ZO methods by estimating gradients using function values. An intuitive approach for estimating the full gradient involves computing partial derivatives along coordinate directions using finite differences. Given a differentiable function $h(\mathbf{z}):\mathbb{R}^{d}\to\mathbb{R}$, the partial gradient along the $i$th coordinate direction is estimated using the finite difference of two function values:
$G_{h,i}(\mathbf{z};r)=\frac{h(\mathbf{z}+r\cdot\mathbf{e}_{i})-h(\mathbf{z})}{r},$ (3)
where $\mathbf{e}_{i}\in\mathbb{R}^{d}$ is the unit vector with only the $i$th entry being 1. The parameter $r>0$, known as the smoothing radius, controls the magnitude of the perturbation to $\mathbf{z}$. By evaluating along each of the $d$ coordinate directions, we obtain the $2d$-point CooGE:
$G_{h}(\mathbf{z};r)=(G_{h,1}(\mathbf{z};r),G_{h,2}(\mathbf{z};r),\cdots,G_{h,d}(\mathbf{z};r))^{T}.$
However, as the dimension $d$ increases, the estimator’s requirement of $2d$ function evaluations may lead to high computational overhead for each single step.
On the other hand, the widely applied $2$-point UniGE estimates a full gradient using only two points [23]:
$G_{h}(\mathbf{z};r;\mathbf{v})=\frac{h(\mathbf{z}+r\cdot\mathbf{v})-h(\mathbf{z})}{r/d}\cdot\mathbf{v},$ (4)
where $\mathbf{v}$ is a perturbation vector uniformly sampled on a unit sphere $\mathcal{U}(\mathbb{S}_{d-1})$.
Finally, we provide two lemmas to characterize the bias and variance of the coordinate-wise partial gradient estimator (3) and $2$-point UniGE (4) under proper assumptions.
Lemma 2.1.
Assume a continuously differentiable function $h$ is $L$-smooth, then, the bias of (3) satisfies:
$\left|G_{h,i}(\mathbf{z};r)-\nabla_{i}h(\mathbf{z})\right|\leq\frac{1}{2}Lr.$ (5)
Lemma 2.2.
Assume a continuously differentiable function $h$ is $G$-Lipschitz and $L$-smooth. Define the smoothed function as $h^{r}(\mathbf{z})=\mathbb{E}_{\bar{\mathbf{v}}\sim\mathcal{U}(\mathbb{B}_{d})}\left[h(\mathbf{z}+r\bar{\mathbf{v}})\right]$. Then, the estimator (4): satisfies:
$\nabla h^{r}(\mathbf{z})=\mathbb{E}_{\mathbf{v}}\left[G_{h}(\mathbf{z};r;\mathbf{v})\right],$ (6)
$\|\nabla h^{r}(\mathbf{z})-\nabla h(\mathbf{z})\|\leq Lr,$ (7)
Furthermore, the variance of (4) is bounded by:
$\mathbb{E}_{\mathbf{v}}\left[\|G_{h}(\mathbf{z};r;\mathbf{v})-\nabla h^{r}(\mathbf{z})\|^{2}\right]\leq\frac{9d^{2}G^{2}}{d+2}+\frac{3d^{2}L^{2}r^{2}}{4}.$ (8)
The proof of Lemma 2.1 is deferred to Appendix A. Lemma 2.2 follows from the continuity and smoothness of $h$, with (6)-(7) provided in [23] and (8) provided in [2]. Lemmas 2.1 and 2.2 suggest that the biases of both estimators can be small by adopting a small $r$, but the coordinate-wise partial gradient estimator enjoys a better variance bound of $\frac{L^{2}r^{2}}{4}$ when $r$ is small.
3   Zeroth-order extra-gradient
In this section, we leverage the $2$-point UniGE to design the zeroth-order extra gradient (ZOEG) algorithm to solve problem (2), and provide its performance guarantees.
3.1   Algorithm design
To solve the min-max problem (2), we first propose the ZOEG algorithm based on (4) and the EG method, as detailed in Algorithm 1. The iterations in our algorithm follow the main steps of extra-gradient as follows:
$\mathbf{z}^{+}_{k}=\mathcal{P}_{\mathcal{Z}}\left[\mathbf{z}_{k}-\eta F(\mathbf{z}_{k})\right],$
$\mathbf{z}_{k+1}=\mathcal{P}_{\mathcal{Z}}\left[\mathbf{z}_{k}-\eta F(\mathbf{z}_{k}^{+})\right],$
where $\eta>0$ is the step size and $\mathbf{z}^{+}_{k}$ is the mid-point to approximate the proximal point (PP) update. Recall that the PP update: $\mathbf{z}_{k+1}=\mathcal{P}_{\mathcal{Z}}\left[\mathbf{z}_{k}-\eta F(\mathbf{z}_{k+1})\right]$ relies on gradient evaluation at $\mathbf{z}_{k+1}$[27]. The implicit nature of the PP update makes it computationally intractable in practice. Therefore, the EG algorithm uses the gradient information at the intermediate point $\mathbf{z}_{k}^{+}$ to approximate the gradient at $\mathbf{z}_{k+1}$, which effectively generates a proximal direction [24].
Given the absence of derivative information, for the $k$th iterate, we sample $\mathbf{v}$ and $\mathbf{w}$ uniformly from the unit sphere $\mathbb{S}_{d-1}$ to construct $\mathbf{g}_{k}$ and $\mathbf{g}_{k}^{+}$ as the estimations of $F(\mathbf{z}_{k})$ and $F(\mathbf{z}_{k}^{+})$ respectively:
$\mathbf{g}_{k}=\frac{f(\mathbf{z}_{k}+r_{k}\mathbf{v}_{k})-f(\mathbf{z}_{k})}{r_{k}/d}\tilde{\mathbf{v}}_{k},$
$\mathbf{g}_{k}^{+}=\frac{f(\mathbf{z}_{k}^{+}+r_{k}\mathbf{w}_{k}-f(\mathbf{z}_{k}^{+})}{r_{k}/d}\tilde{\mathbf{w}}_{k},$
where $\tilde{\mathbf{v}}_{k}=[\mathbf{v}_{k}(1:d_{x});-\mathbf{v}_{k}(d_{x}+1:d)]$, and $\tilde{\mathbf{w}}_{k}=[\mathbf{w}_{k}(1:d_{x});-\mathbf{w}_{k}(d_{x}+1:d)]$. According to Lemma 2.2, it can be shown that they are good estimators of $F(\mathbf{z}_{k})$ and $F(\mathbf{z}_{k}^{+})$ with bounded bias under Assumption 2.2, so we replace the first-order derivatives in extra-gradient method with $\mathbf{g}_{k}$ and $\mathbf{g}_{k}^{+}$ respectively [23].
Note that we sample $\mathbf{v}_{k}$ (and $\mathbf{w}_{k}$) from $\mathcal{U}(\mathbb{S}_{d-1})$ to estimate $F(\mathbf{z}_{k})$ (and $F(\mathbf{z}_{k}^{+})$) jointly. we can also estimate the partial gradients in $\mathbf{x}$ and $\mathbf{y}$ separately by adding two perturbations from $\mathcal{U}(\mathbb{S}_{d_{x}-1})$ and $\mathcal{U}(\mathbb{S}_{d_{y}-1})$ to $\mathbf{x}$ and $\mathbf{y}$ respectively. Moreover, the ZOEG algorithm outputs the averaged value over all iterations, which is only for the convenience of theoretical analysis. In practice, we can take the last iterate as the final decision.
Algorithm 1 Zeroth-Order Extra-Gradient (ZOEG)
  Inputs: Number of iterations $K$, step sizes $\eta$, smoothing radii $\{r_{k}\}_{k=0}^{K-1}$, and initial point $\mathbf{z}_{0}$.
  for $k\xleftarrow{}0$ to $K-1$ do
     Sample $\mathbf{v}_{k}$ and $\mathbf{w}_{k}$ from $\mathcal{U}(\mathbb{S}_{d-1})$ independently.
     Update $\mathbf{z}_{k}$ by:
$\mathbf{z}^{+}_{k}=\mathcal{P}_{\mathcal{Z}}\left[\mathbf{z}_{k}-\eta\mathbf{g}_{k}\right],$ (9)
$\mathbf{z}_{k+1}=\mathcal{P}_{\mathcal{Z}}\left[\mathbf{z}_{k}-\eta\mathbf{g}_{k}^{+}\right].$ (10)
  end for
  Outputs: $\hat{\mathbf{z}}_{K}=\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{z}^{+}_{k}.$
3.2   Finite-sample analysis
The following theorem provides the convergence results for the proposed ZOEG algorithm.
Theorem 3.1.
Suppose Assumptions 2.1-2.3 hold. The sequences $\{\mathbf{z}_{k}\}$, and $\{\mathbf{z}_{k}^{+}\}$ are the iterates generated by ZOEG Algorithm. Set the smoothing radii to satisfy $\sum_{k=1}^{\infty}r_{k}=M_{1}<\infty$ and $\sum_{k=1}^{\infty}r_{k}^{2}=M_{2}<\infty$. Then for all $K>1$, the duality gap of the averaged iterates $\hat{\mathbf{z}}_{K}=\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{z}_{k}$ satisfies
$\mathbb{E}\left[\Delta(\hat{\mathbf{z}}_{K})\right]\leq\frac{\tilde{D}^{2}}{2\eta K}+\frac{C_{1}\eta+C_{2}}{K}+108\eta dG^{2},$
where $C_{1}=(9L^{2}+12)d^{2}M_{2}$, $C_{2}=L\tilde{D}M_{1}$.
Proof of Theorem 3.1.  We start by defining the following filtration $\mathcal{F}_{k}:=\sigma(\mathbf{z}_{0},\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}_{1},\mathbf{v}_{1},\mathbf{w}_{1},...,\mathbf{z}_{k}).$ Then, the upper bound on the conditional expectation of zeroth-order EG iterates $\left\langle\eta F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle$ can be established in the following lemma, the proof of which is given in Appendix A.1.
Lemma 3.1.
Under Assumption 2.12.3, given the zeroth-order extra-gradient update rules (9), (10), we have for any $\mathbf{z}\in\mathcal{Z}$:
$\ \mathbb{E}\left[\left\langle\eta F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\mid\mathcal{F}_{k}\right]\leq\mathbb{E}\left[H_{k}(\mathbf{z})-H_{k+1}(\mathbf{z})+\Delta_{k}\mid\mathcal{F}_{k}\right],$ (11)
where $H_{k}(\mathbf{z})$ and $\Delta_{k}$ are defined as $H_{k}(\mathbf{z})=\frac{1}{2}\|\mathbf{z}_{k}-\mathbf{z}\|^{2},\Delta_{k}=\!\left(\frac{108d^{2}}{d+2}+12\right)\eta^{2}G^{2}\!+\eta^{2}L^{2}(9d^{2}\!+\!12)r_{k}^{2}\!+\eta L\tilde{D}r_{k}.$
Given Lemma 3.1, taking the total expectation of (11) leads to:
$\mathbb{E}\left[\left\langle\eta F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]\leq\mathbb{E}\left[H_{k}(\mathbf{z})-H_{k+1}(\mathbf{z})\right]+\Delta_{k}.$
Then, by calculating the telescoping sum of $\mathbb{E}[\langle F(\mathbf{z}_{k}^{+}),\allowbreak\mathbf{z}_{k}^{+}-\mathbf{z}\rangle]$ for $k=0,1,\cdots,K-1$, we can bound $\sum_{k=0}^{K-1}\mathbb{E}[\langle F(\mathbf{z}_{k}^{+}),\allowbreak\mathbf{z}_{k}^{+}-\mathbf{z}\rangle]$ as follows:
$\ \sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]$
$\leq$ $\ \sum_{k=0}^{K-1}\mathbb{E}\left[\frac{H_{k}(\mathbf{z})}{\eta}-\frac{H_{k+1}(\mathbf{z})}{\eta}\right]+\sum_{k=0}^{K-1}\frac{\Delta_{k}}{\eta}$
$=$ $\ \mathbb{E}\left[\frac{H_{0}(\mathbf{z})}{\eta}-\frac{H_{K}(\mathbf{z})}{\eta}\right]+\sum_{k=0}^{K-1}\frac{\Delta_{k}}{\eta}$
$\leq$ $\ \frac{1}{\eta}\mathbb{E}\left[\frac{1}{2}\|\mathbf{z}_{0}-\mathbf{z}\|^{2}\right]+\sum_{k=0}^{K-1}\frac{\Delta_{k}}{\eta}$
$\leq$ $\ \frac{\tilde{D}^{2}}{2\eta}+\sum_{k=0}^{K-1}\frac{\Delta_{k}}{\eta},$
where the second inequality follows from the definition of $H_{k}(\mathbf{z})$ and the fact $\frac{1}{2}\|\mathbf{z}_{K}-\mathbf{z}\|^{2}\geq 0$, and the last inequality comes from the boundedness of $\mathcal{Z}$. Consequently, we have:
$\sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]\leq\frac{\tilde{D}^{2}}{2\eta}+C_{1}\eta+C_{2}+108K\eta dG^{2}.$ (12)
Then, the relation between the telescoping sum $\sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]$ and the duality gap is given in the following lemma.
Lemma 3.2.
Suppose Assumptions 2.1-2.3 hold. When there exists some bound $\Gamma_{K}\geq 0$ satisfying $\sum_{k=0}^{K-1}\mathbb{E}[\langle F(\mathbf{z}_{k}^{+}),\allowbreak\mathbf{z}_{k}^{+}-\mathbf{z}\rangle]\leq\Gamma_{K}$ for all $K>1$, $\mathbf{z}\in\mathcal{Z}$, the duality gap of the averaged iterate $\hat{\mathbf{z}}_{K}$ is bounded as
$\mathbb{E}\left[\Delta(\hat{\mathbf{z}}_{K})\right]\leq\frac{\Gamma_{K}}{K}.$
Proof of Lemma 3.2.  Following the definition of $(\hat{\mathbf{x}}_{K},\hat{\mathbf{y}}_{K})$, the duality gap $\mathbb{E}\left[f\left(\hat{\mathbf{x}}_{K},\mathbf{y}\right)-f\left(\mathbf{x},\hat{\mathbf{y}}_{K}\right)\right]$ is upper bounded by:
$\mathbb{E}\left[f\left(\hat{\mathbf{x}}_{K},\mathbf{y}^{*}\right)-f\left(\mathbf{x}^{*},\hat{\mathbf{y}}_{K}\right)\right]$
$=$ $\mathbb{E}\left[f\left(\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{x}_{k}^{+},\mathbf{y}^{*}\right)-f\left(\mathbf{x}^{*},\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{y}_{k}^{+}\right)\right]$
$\leq$ $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[f(\mathbf{x}_{k}^{+},\mathbf{y}^{*})-f(\mathbf{x}^{*},\mathbf{y}_{k}^{+})\right]$
$=$ $\frac{1}{K}\!\sum_{k=0}^{K-1}\!\mathbb{E}\left[f(\mathbf{x}_{k}^{+},\mathbf{y}^{*})\!-\!f(\mathbf{z}_{k}^{+})\!+\!f(\mathbf{z}_{k}^{+})\!-\!f(\mathbf{x}^{*},\mathbf{y}_{k}^{+})\right]$
$\leq$ $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}^{*}\right\rangle\right]$
$\leq$ $\frac{\Gamma_{K}}{K},$
where the first inequality follows from Jensen’s inequality and the second one follows from that $f$ is convex in $\mathbf{x}$ and concave in $\mathbf{y}$.
Finally, combining the results in (12) and Lemma 3.2 finishes the proof.
Theorem 3.1 shows that, under constant step sizes, the expected duality gap decays at the rate of $\mathcal{O}(1/K)$ up to a fixed error, which can approach zero by employing diminishing or sufficiently small step sizes.
The theoretical guarantee in Theorem 3.1 only characterizes the finite-iteration bound of the duality gap of problem (2), while the behavior of the averaged iterate $\hat{\mathbf{x}}_{k}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{x}_{k}^{+}$ w.r.t the original constrained problem (1) remains unclear. To bridge this gap, we present the following proposition, which gives explicit bounds on the optimality gap and constraint violation of $\hat{\mathbf{z}}_{K}$ for problem (1).
Proposition 3.1.
Suppose the assumptions and conditions in Theorem 3.1 hold. Then for any $K>1$, we have
$\|\left[\mathbb{E}\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]\right]_{+}\|$ $\leq\frac{\tilde{D}^{2}}{2\eta RK}+\frac{C_{1}\eta+C_{2}}{RK}+\frac{108\eta dG^{2}}{R},$
$\mathbb{E}\left[\phi_{0}(\hat{\mathbf{x}}_{K})\right]-\phi_{0}^{*}$ $\leq\frac{\tilde{D}^{2}}{2\eta K}+\frac{C_{1}\eta+C_{2}}{K}+108\eta dG^{2}.$
Proof of Proposition 3.1.  Recall that Lemma 3.2 establishes the upper bound for the duality gap in solving the min-max formulation. Following a similar idea, we provide the following lemma to link the telescoping sum $\sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]$ to both the performance in the optimality and constaint violation of the derived solution $\hat{\mathbf{x}}_{K}$.
Lemma 3.3.
Suppose Assumptions 2.1-2.3 hold. When there exists some bound $\Gamma_{K}\geq 0$ satisfying $\sum_{k=0}^{K-1}\mathbb{E}[\langle F(\mathbf{z}_{k}^{+}),\allowbreak\mathbf{z}_{k}^{+}-\mathbf{z}\rangle]\leq\Gamma_{K}$ for any $K>1$, and $\mathbf{z}\in\mathcal{Z}$, the expected optimality gap and constraint violation are bounded by
$\left\|\left[\mathbb{E}\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]\right]_{+}\right\|$ $\leq\frac{\Gamma_{K}}{RK},$
$\mathbb{E}\left[\phi_{0}(\hat{\mathbf{x}}_{K})\right]-\phi_{0}^{*}$ $\leq\frac{\Gamma_{K}}{K},$
where $R=\max\{r>0|\mathbf{y}^{*}+r\mathbf{u}\in\mathcal{Y},\|\mathbf{u}\|=1,\forall\mathbf{u}\in\mathbb{R}^{d_{y}}\}$.
Proof of Lemma 3.3.  We start from the upper bound on $\left\langle\mathbf{y}-\mathbf{y}^{*},\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle,\forall\mathbf{y}\in\mathcal{Y}$, as follows:
$\ \left\langle\mathbf{y}-\mathbf{y}^{*},\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle$
$=$ $\ \left\langle\mathbf{y}-\mathbf{y}_{k}^{+},\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle+\left\langle\mathbf{y}_{k}^{+}-\mathbf{y}^{*},\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle$
$\leq$ $\ \left\langle\mathbf{y}_{k}^{+}-\mathbf{y},-\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle+f(\mathbf{z}_{k}^{+})-f(\mathbf{x}_{k}^{+},\mathbf{y}^{*})$
$\leq$ $\ \left\langle\mathbf{y}_{k}^{+}-\mathbf{y},-\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle-f(\mathbf{x}_{k}^{+},\mathbf{y}^{*})+\left\langle\mathbf{x}_{k}^{+}-\mathbf{x}^{*},\nabla f_{\mathbf{x}}(\mathbf{z}_{k}^{+})\right\rangle+f(\mathbf{x}^{*},\mathbf{y}_{k}^{+})$
$\leq$ $\ \left\langle\mathbf{y}_{k}^{+}-\mathbf{y},-\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle+\left\langle\mathbf{x}_{k}^{+}-\mathbf{x}^{*},\nabla f_{\mathbf{x}}(\mathbf{z}_{k}^{+})\right\rangle,$ (13)
where the first inequality follows from the concavity of $f$ in $\mathbf{y}$, the second inequality follows from the convexity of $f$ in $\mathbf{x}$, and the last inequality follows from that $\mathbf{z}^{*}$ is the saddle point, which indicates $f(\mathbf{x}^{*},\mathbf{y}_{k}^{+})\leq f(\mathbf{x}^{*},\mathbf{y}^{*})\leq f(\mathbf{x}_{k}^{+},\mathbf{y}^{*})$.
Denote $\tilde{\mathbf{z}}:=(\mathbf{x}^{*},\mathbf{y})$, taking the total expectation of (3.2) and then taking the telescoping sum lead to:
$\ \mathbb{E}\left[\left\langle\mathbf{y}-\mathbf{y}^{*},\sum_{k=0}^{K-1}\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle\right]$
$\leq$ $\ \sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle\mathbf{y}_{k}^{+}-\mathbf{y},-\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle+\left\langle\mathbf{x}_{k}^{+}-\mathbf{x}^{*},\nabla f_{\mathbf{x}}(\mathbf{z}_{k}^{+})\right\rangle\right]$
$=$ $\ \sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle\mathbf{z}_{k}^{+}-\tilde{\mathbf{z}},F(\mathbf{z}_{k}^{+})\right\rangle\right]$
$\leq$ $\ \Gamma_{K}.$
Define $\mathbf{s}:=\mathbb{E}\left[\sum_{k=0}^{K-1}\nabla f_{\mathbf{y}}(\mathbf{x}_{k}^{+},\mathbf{y}_{k}^{+})\right]$. By the convexity of $f$ in $\mathbf{x}$, we have $\mathbf{s}\geq K\mathbb{E}\left[\nabla f_{\mathbf{y}}(\hat{\mathbf{x}}_{K},\hat{\mathbf{y}}_{K})\right]=K\mathbb{E}\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]$. If $\left[\mathbf{s}\right]_{+}=0$, we have $[\mathbb{E}\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]]_{+}=0$ and the constraint violation bound holds. Otherwise, define $\bar{\mathbf{y}}=\mathbf{y}^{*}+R\frac{\left[\mathbf{s}\right]_{+}}{\|\left[\mathbf{s}\right]_{+}\|}$. By the definition of $R$, $\bar{\mathbf{y}}$ lies within $\mathcal{Y}$. Given that $\left\langle\bar{\mathbf{y}}-\mathbf{y}^{*},\mathbb{E}\left[\sum_{k=0}^{K-1}\nabla f_{\mathbf{y}}(\mathbf{x}_{k}^{+},\mathbf{y}_{k}^{+})\right]\right\rangle=R\|\left[\mathbf{s}\right]_{+}\|\leq\Gamma_{K}$, we have:
$\|\mathbb{E}\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]_{+}\|\leq\frac{\Gamma_{K}}{RK}.$
As for the optimal gap of the objective function value, we apply the convexity of the objective function $\phi_{0}$ in $\mathbf{x}$ to get:
$\ \phi_{0}(\hat{\mathbf{x}}_{K})-\phi_{0}^{*}$
$\leq$ $\ \frac{1}{K}\sum_{k=0}^{K-1}\left[\phi_{0}(\mathbf{x}_{k}^{+})-\phi_{0}^{*}\right]$
$=$ $\ \frac{1}{K}\sum_{k=0}^{K-1}\left[f(\mathbf{x}_{k}^{+},\mathbf{y}^{*})-\left\langle\mathbf{y}^{*},\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle-\phi_{0}^{*}\right]$
$\leq$ $\ \frac{1}{K}\sum_{k=0}^{K-1}\left[f(\mathbf{x}_{k}^{+},\mathbf{y}_{k}^{+})+\left\langle\mathbf{y}_{k}^{+},-\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle-\phi_{0}^{*}\right]$
$\leq$ $\ \frac{1}{K}\sum_{k=0}^{K-1}\left[\left\langle\mathbf{x}_{k}^{+}-\mathbf{x}^{*},\nabla f_{\mathbf{x}}(\mathbf{z}_{k}^{+})\right\rangle+f(\mathbf{x}^{*},\mathbf{y}_{k}^{+})\right]+\frac{1}{K}\sum_{k=0}^{K-1}\left[\left\langle\mathbf{y}_{k}^{+},-\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle-f(\mathbf{x}^{*},\mathbf{y}^{*})\right]$
$\leq$ $\ \frac{1}{K}\sum_{k=0}^{K-1}\left[\left\langle\mathbf{x}_{k}^{+}-\mathbf{x}^{*},\nabla f_{\mathbf{x}}(\mathbf{z}_{k}^{+})\right\rangle\right]+\frac{1}{K}\sum_{k=0}^{K-1}\left[\left\langle\mathbf{y}_{k}^{+},-\nabla f_{\mathbf{y}}(\mathbf{z}_{k}^{+})\right\rangle\right],$
where the first inequality follows from Jensen’s inequality; the second inequality follows from the concavity of $f$ in $\mathbf{y}$; the third inequality follows from the convexity of $f$ in $\mathbf{x}$; and the last one is derived from $f(\mathbf{x}^{*},\mathbf{y}_{k}^{+})\leq f(\mathbf{x}^{*},\mathbf{y}^{*})$. Let $\check{\mathbf{z}}:=(\mathbf{x}^{*},\mathbf{0})\in\mathcal{Z}$, then we have:
$\sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\check{\mathbf{z}}\right\rangle\right]\leq\Gamma_{K}.$
Consequently, the optimal gap of $\phi_{0}(\hat{\mathbf{x}}_{K})$ can be bounded by $\mathbb{E}\left[\phi_{0}(\hat{\mathbf{x}}_{K})\right]-\phi_{0}^{*}\leq\frac{\Gamma_{K}}{K}$.
Substituting (12) into Lemma 3.3 finishes the proof.
It is shown in Proposition 3.1 that the constraint violation and the optimality gap can both converge at the same convergence rate with the duality gap. For the bounds in Theorem 3.1 and Proposition 3.1, we can apply small step sizes to eliminate the fixed error terms. As a corollary, the iteration complexity required to obtain an $\epsilon$-saddle point for problem (2) and an $\epsilon$-optimal solution for problem (1) is provided as follows.
Corollary 3.2.
Suppose the assumptions and conditions in Theorem 3.1 hold. Set $\eta=\frac{\tilde{D}}{6\sqrt{6dK}G}$ and $K\geq\max\left\{\frac{864dG^{2}\tilde{D}^{2}}{\epsilon^{2}},\left(\frac{(6d^{2}+8)^{2}L^{4}M_{2}^{2}\tilde{D}^{2})}{6dG^{2}\epsilon^{2}}\right)^{\frac{1}{3}},\frac{4LM_{1}\tilde{D}}{\epsilon}\right\}$ Then, we have:
$\mathbb{E}\left[\Delta(\hat{\mathbf{z}}_{K})\right]\leq\epsilon,$
$\|\left[\mathbb{E}\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]\right]_{+}\|\leq\frac{\epsilon}{R},$
$\mathbb{E}\left[\phi_{0}(\hat{\mathbf{x}}_{K})\right]-\phi_{0}^{*}\leq\epsilon.$
Proof of Corollary 3.2.  By the results in Theorem 3.1, substituting $\eta=\frac{\tilde{D}}{6\sqrt{6dK}G}$ into $\frac{\frac{\tilde{D}^{2}}{2\eta}+C_{1}\eta+C_{2}}{K}+108\eta dG^{2}$ gives:
$\ \mathbb{E}\left[\Delta(\hat{\mathbf{z}}_{K})\right]$
$\leq$ $\ \frac{\tilde{D}^{2}}{2\eta K}+108\eta dG^{2}+\frac{(9d^{2}+12)\eta L^{2}M_{2}+L\tilde{D}M_{1}}{K}$
$=$ $\ 6\sqrt{\frac{6d}{K}}\tilde{D}G+\frac{(3d^{2}+4)L^{2}M_{2}\tilde{D}}{2\sqrt{6dK^{3}}G}+\frac{LM_{1}\tilde{D}}{K}$
$\leq$ $\ \frac{\epsilon}{2}+\frac{\epsilon}{4}+\frac{\epsilon}{4}$
$=$ $\ \epsilon,$
where the first equality follows from $\eta=\frac{\tilde{D}}{6\sqrt{6dK}G}$ and the last inequality follows from the lower bound of $K$. The remaining two bounds follow a similar proof.
The above corollary indicates that the iteration complexity to achieve an $\epsilon$-approximate solution for problems (1) and (2) is $\mathcal{O}\left(\frac{d}{\epsilon^{2}}\right)$, which is also the oracle complexity. The dependence on $\epsilon$ matches the best-known result of $\mathcal{O}(\frac{d^{2}}{\epsilon^{2}})$ in [30] for solving problem (1) and another result of $\mathcal{O}(\frac{d^{4}}{\epsilon^{2}})$ in [22] for solving problem (2). Moreover, our results achieve a better and the best-known dependence on $d$.
Remark.
While our analysis focuses on the deterministic setting, ZOEG and the main theoretical results are also applicable to the stochastic case, as discussed in [30]. Stochasticity introduces an additional variance term in the gradient estimation bound (see (8)), but our complexity results remain unchanged.
Recall that the oracle complexity of the first-order EG is $\mathcal{O}(\frac{1}{\epsilon})$ [27], which means ZOEG is strictly inferior to the first-order method. Therefore, it is very natural to ask whether we can improve ZOEG to bridge the gap. In the next section, we provide a positive answer to this question by leveraging the $2d$-point coordinate gradient estimator.
4   Zeroth-order coordinate extra-gradient
In this section, we will present another algorithm called zeroth-order coordinate extra-gradient (ZOCEG) and provide its convergence guarantees.
4.1   Algorithm design
According to the analysis in Section 3, ZOEG suffers from slow convergence. The main factor that limits the performance of ZOEG is the large variance of $2$-point UniGE (4). By Lemma 2.1, the coordinate-wise partial gradient estimator (3) enjoys controllable bias and variance by using small smoothing radii. To leverage such merits, we integrate $2d$-point CooGE into the EG method and develop the ZOCEG algorithm. The gradient estimators of $\nabla_{\mathbf{x}}f(\mathbf{x}_{k},\mathbf{y}_{k})$ and $\nabla_{\mathbf{y}}f(\mathbf{x}_{k},\mathbf{y}_{k})$ are given by:
$\hat{\mathbf{g}}^{x}_{k}=\sum_{i=1}^{d_{x}}\frac{f(\mathbf{x}_{k}+r_{k}\mathbf{e}_{i},\mathbf{y}_{k})-f(\mathbf{x}_{k},\mathbf{y}_{k})}{r_{k}}\mathbf{e}_{i},$
$\hat{\mathbf{g}}^{y}_{k}=\sum_{j=1}^{d_{y}}\frac{f(\mathbf{x}_{k},\mathbf{y}_{k}+r_{k}\mathbf{e}_{j})-f(\mathbf{x}_{k},\mathbf{y}_{k})}{r_{k}}\mathbf{e}_{j}.$
With $\mathbf{x}_{k}$ and $\mathbf{y}_{k}$ updated by the estimated gradients, we get $(\mathbf{x}_{k}^{+},\mathbf{y}_{k}^{+})$, and the gradient of $f(\mathbf{x}_{k}^{+},\mathbf{y}_{k}^{+})$ can be estimated similarly as follows:
$\hat{\mathbf{g}}^{x,+}_{k}=\sum_{i=1}^{d_{x}}\frac{f(\mathbf{x}^{+}_{k}+r_{k}\mathbf{e}_{i},\mathbf{y}^{+}_{k})-f(\mathbf{x}^{+}_{k},\mathbf{y}^{+}_{k})}{r_{k}}\mathbf{e}_{i},$
$\hat{\mathbf{g}}^{y,+}_{k}=\sum_{j=1}^{d_{y}}\frac{f(\mathbf{x}^{+}_{k},\mathbf{y}^{+}_{k}+r_{k}\mathbf{e}_{j})-f(\mathbf{x}^{+}_{k},\mathbf{y}^{+}_{k})}{r_{k}}\mathbf{e}_{j}.$
The detailed procedure is presented in Algorithm 2.
Algorithm 2 Zeroth-Order Coordinate Extra-Gradient (ZOCEG)
  Inputs: Number of iterations $K$, step size $\eta$, smoothing radii $\{r_{k}\}$, and initial point $\mathbf{z}_{0}$.
  for $k\xleftarrow{}0$ to $K-1$ do
     Calculate $\mathbf{z}_{k}^{+}$ by:
$\mathbf{x}^{+}_{k}=\mathcal{P}_{\mathcal{X}}\left[\mathbf{x}_{k}-\eta\hat{\mathbf{g}}_{k}^{x}\right],$
$\mathbf{y}^{+}_{k}=\mathcal{P}_{\mathcal{Y}}\left[\mathbf{y}_{k}+\eta\hat{\mathbf{g}}_{k}^{y}\right].$
Based on $\mathbf{z}_{k}^{+}$, update $\mathbf{z}_{k}$ by:
$\mathbf{x}_{k+1}=\mathcal{P}_{\mathcal{X}}\left[\mathbf{x}_{k}-\eta\hat{\mathbf{g}}_{k}^{x,+}\right],$
$\mathbf{y}_{k+1}=\mathcal{P}_{\mathcal{Y}}\left[\mathbf{y}_{k}+\eta\hat{\mathbf{g}}_{k}^{y,+}\right].$
  end for
  Outputs: $\hat{\mathbf{z}}_{K}=\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{z}^{+}_{k}.$
4.2   Finite-sample analysis
First, we focus on the performance of ZOCEG when solving problem (2). The duality gap of the averaged iterate from ZOCEG is bounded in the following theorem.
Theorem 4.1.
Suppose Assumptions 2.1-2.3 hold. The sequence $\{\mathbf{z}_{k}^{+}\}$ is generated by the ZOCEG Algorithm. Set the smoothing radii to satisfy $\sum_{k=1}^{\infty}r_{k}\leq\frac{M_{3}}{\sqrt{d_{x}}+\sqrt{d_{y}}}<\infty$ for some $M_{3}>0$. Let the step sizes satisfy $0<\eta L\leq\frac{1}{2}$, then for any $K>1$, we have:
$\Delta(\hat{\mathbf{z}}_{K})\leq\frac{1}{2K}\left(\frac{\tilde{D}^{2}}{\eta}+3LM_{3}\tilde{D}\right).$
Proof of Theorem 4.1.  Following a similar idea to analyzing ZOEG, we start from providing the bound of $\left\langle\eta F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle$ in the following lemma, the proof of which is deferred to Appendix B.
Lemma 4.1.
Suppose Assumptions 2.1-2.3 hold and the step sizes satisfy $0<\eta L\leq\frac{1}{2}$. Given the iterates from ZOCEG, we have for any $\mathbf{z}\in\mathcal{Z}$:
$\left\langle\eta F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\leq H_{k}(\mathbf{z})-H_{k+1}(\mathbf{z})+\frac{3}{2}\eta Lr_{k}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}.$
Based on Lemma 4.1, taking the telescoping sum of $\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle$ gives the upper bound as follows:
$\ \sum_{k=0}^{K-1}\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle$
$\leq$ $\ \frac{1}{\eta}H_{0}(\mathbf{z})-\frac{1}{\eta}H_{K}(\mathbf{z})+\sum_{k=0}^{K-1}\frac{3}{2}Lr_{k}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}$
$\leq$ $\ \frac{\tilde{D}^{2}}{2\eta}+\frac{3}{2}L\frac{M_{3}}{\sqrt{d_{x}}+\sqrt{d_{y}}}\cdot(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}$
$\leq$ $\frac{\tilde{D}^{2}}{2\eta}+\frac{3}{2}LM_{3}\tilde{D},$ (14)
where the last inequality follows from the definition of $H_{k}(\mathbf{z})$ and the boundedness of $\mathcal{Z}$. Substituting (4.2) into Lemma 3.2, which still holds without the expectation, finishes the proof.
The result in Theorem 4.1 shows that the generated iterates from ZOCEG can converge to a saddle point at the convergence rate of $\mathcal{O}\left(\frac{1}{K}\right)$. Following the same idea in the analysis of ZOEG, combining (4.2) with the deterministic counterpart of Lemma 3.3 directly establishes the following proposition, which provides finite-iteration bounds on the optimality gap and constraint violation of ZOCEG for solving the original constrained problem (1). Therefore, we omit its proof here.
Proposition 4.1.
Suppose the Assumptions and conditions in Theorem 4.1 hold. Then for any $K>1$, we have:
$\|\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]_{+}\|$ $\leq\frac{\tilde{D}^{2}}{2RK\eta}+\frac{3LM_{3}\tilde{D}}{2RK},$
$\phi_{0}(\hat{\mathbf{x}}_{K})-\phi_{0}^{*}$ $\leq\frac{\tilde{D}^{2}}{2K\eta}+\frac{3LM_{3}\tilde{D}}{2K}.$
Based on Theorem 4.1 and Proposition 4.1, we can further establish the oracle complexity of ZOCEG when solving problems (1) and (2).
Corollary 4.2.
Suppose the Assumptions and conditions in Theorem 4.1 hold. Then for any $K\geq\max\left\{\frac{\tilde{D}^{2}}{\eta\epsilon},\frac{3LM_{3}\tilde{D}}{\epsilon}\right\}$, we have
$\Delta(\hat{\mathbf{z}}_{K})\leq\epsilon,$
$\|\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]^{+}\|\leq\frac{\epsilon}{R},$
$\mathbb{E}\left[\phi_{0}(\hat{\mathbf{x}}_{K})\right]-\phi_{0}^{*}\leq\epsilon.$
The proof of Corollary 4.2 follows the same idea as Corollary 3.2 and is thus omitted. Corollary 4.2 shows that the oracle complexity for ZOCEG to derive an $\epsilon$-saddle point for problem (2) and an $\epsilon$-optimal solution for problem (1) is $\mathcal{O}\left(\frac{d}{\epsilon}\right)$, which differs from first-order EG by a factor of $d$ [27]. This is the best-known complexity bound for zeroth-order algorithms when solving problem (1).
Remark.
It is worth noting that the complexity bounds proportional to $\epsilon^{-1}$ are also derived in [14] and [17] for deterministic constrained ZO algorithms. However, both studies rely on a decomposable and explicit feasible set, which cannot apply to problem (1) with black-box constraints.
Remark.
Compared to ZOEG, the oracle complexity is improved from $\mathcal{O}\left(\frac{d}{\epsilon^{2}}\right)$ to $\mathcal{O}\left(\frac{d}{\epsilon}\right)$. The key factor driving the improvement is that the bias and variance of the $2d$-point CooGE can be controlled to be sufficiently small, which makes ZOCEG resemble the behavior of first-order EG. However, ZOCEG and its theoretical results do not apply to the stochastic setting, where extra variance is introduced to gradient estimation due to the system’s stochasticity.
Remark.
We consider the convex problems throughout this work. As for nonconvex problems, we can combine our algorithms with the PP method in [30] and [3], a meta-algorithm, to design a double-loop algorithm. Using the theoretical results in [3], it is not hard to derive the oracle complexities of $\mathcal{O}\left(\frac{d}{\epsilon^{3}}\right)$ and $\mathcal{O}\left(\frac{d}{\epsilon^{2}}\right)$ for ZOEG and ZOCEG-based PP algorithms to get an $\epsilon$-critical KKT point in nonconvex problems. The result of $\mathcal{O}\left(\frac{d}{\epsilon^{2}}\right)$ also matches recent studies for nonconvex ZO problems [13, 39]. When directly applying our algorithms to nonconvex problems, we leave the complexity analysis for future work and only conduct empirical validation in this study.
4.3   Zeroth-order block coordinate extra-gradient
While ZOCEG has excellent oracle complexity guarantees, each gradient estimation requires $\mathcal{O}(d)$ function values, which is inefficient for high-dimensional problems. To improve computational efficiency for each single step, we may further consider only estimating partial gradients along a block of coordinates rather than the full gradient and updating the block of coordinates accordingly. Let $\tau_{x},\tau_{y}$ denote the number of coordinates updated for the primal and dual variables at each iteration, where $1\leq\tau_{x}\leq d_{x}$, and $1\leq\tau_{y}\leq d_{y}$. At the $k$th iteration, the coordinate index subsets $\mathcal{I}_{k}^{x}\subseteq\{1,2,\cdots,d_{x}\}$ and $\mathcal{I}_{k}^{y}\subseteq\{1,2,\cdots,d_{y}\}$, with $|\mathcal{I}_{k}^{x}|=\tau_{x}$, $|\mathcal{I}_{k}^{y}|=\tau_{y}$, are sampled uniformly without replacement. The partial gradient vectors $\left\{\nabla_{\mathbf{x},i}f(\mathbf{x}_{k},\mathbf{y}_{k})\right\}_{i\in\mathcal{I}_{k}^{x}}$ and $\left\{\nabla_{\mathbf{y},j}f(\mathbf{x}_{k},\mathbf{y}_{k})\right\}_{j\in\mathcal{I}_{k}^{y}}$ corresponding to selected coordinates can then be estimated as follows:
$\tilde{\mathbf{g}}^{x}_{k}=\sum_{i\in\mathcal{I}_{k}^{x}}\frac{f(\mathbf{\mathbf{x}}_{k}+r_{k}\mathbf{e}_{i},\mathbf{y}_{k})-f(\mathbf{\mathbf{x}}_{k},\mathbf{y}_{k})}{r_{k}}\mathbf{e}_{i},$
$\tilde{\mathbf{g}}^{y}_{k}=\sum_{j\in\mathcal{I}_{k}^{y}}\frac{f(\mathbf{\mathbf{x}}_{k},\mathbf{y}_{k}+r_{k}\mathbf{e}_{j})-f(\mathbf{\mathbf{x}}_{k},\mathbf{y}_{k})}{r_{k}}\mathbf{e}_{j}.$
[…]
$\tilde{\mathbf{g}}^{x,+}_{k}=\sum_{p\in\mathcal{J}_{k}^{x}}\frac{f(\mathbf{\mathbf{x}}^{+}_{k}+r_{k}\mathbf{e}_{p},\mathbf{y}_{k}^{+})-f(\mathbf{\mathbf{x}}_{k}^{+},\mathbf{y}_{k}^{+})}{r_{k}}\mathbf{e}_{p},$
$\tilde{\mathbf{g}}^{y,+}_{k}=\sum_{q\in\mathcal{J}_{k}^{y}}\frac{f(\mathbf{\mathbf{x}}^{+}_{k},\mathbf{y}_{k}^{+}+r_{k}\mathbf{e}_{q})-f(\mathbf{\mathbf{x}}^{+}_{k},\mathbf{y}^{+}_{k})}{r_{k}}\mathbf{e}_{q},$
which are then used to compute the new iterates $\left(\mathbf{x}_{k+1},\mathbf{y}_{k+1}\right)$.
The detailed algorithm is presented in Algorithm 3. Note that the blocks for gradient estimation of $(\mathbf{x}_{k},\mathbf{y}_{k})$ and $(\mathbf{x}_{k}^{+},\mathbf{y}_{k}^{+})$ are sampled independently. The convergence analysis of ZOBCEG is provided in Theorem 4.3, the proof of which is deferred to Appendix C.
Theorem 4.3.
Suppose Assumptions 2.1-2.3 hold. The sequence $\{\mathbf{z}_{k}^{+}\}$ is generated by the ZOBCEG Algorithm. Set the smoothing radii to satisfy $\sum_{k=1}^{\infty}r_{k}=M_{1}<\infty$ and $\sum_{k=1}^{\infty}r_{k}^{2}=M_{2}<\infty$ , then for any $K>1$, we have:
$\mathbb{E}\left[\Delta(\hat{\mathbf{z}}_{K})\right]\leq$ $\ \left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)\frac{\tilde{D}^{2}}{2\eta K}+2\left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)\eta G^{2}+\frac{LM_{1}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2K}+\frac{d_{x}+d_{y}}{2K}\eta L^{2}M_{2}$
$\ +\frac{2\eta GLM_{1}}{K}(\frac{d_{x}}{\sqrt{\tau_{x}}}+\frac{d_{y}}{\sqrt{\tau_{y}}}).$
Based on the result in Theorem 4.3, ZOBCEG can achieve the convergence to a saddle point at the rate of $\mathcal{O}\left(\frac{\tilde{d}}{\eta K}+\frac{\eta d}{K}+\tilde{d}\eta\right)$, where $\tilde{d}=\max\left\{\frac{d_{x}}{\tau_{x}},\frac{d_{y}}{\tau_{y}}\right\}$. Similarly to ZOEG, we can apply a small step size of $\eta$ to achieve the convergence rate of $\mathcal{O}\left(\frac{\tilde{d}}{\sqrt{K}}\right)$. Furthermore, the oracle complexity of ZOBCEG is $\mathcal{O}\left(\frac{d}{\epsilon^{2}}\right)$ to derive an $\epsilon$-saddle point for problem (2) and an $\epsilon$-optimal solution for problem (1).
Compared with ZOCEG, the theoretical guarantees of ZOBCEG degrade due to the random selection of a block of coordinates for updates, which introduces further variance for gradient estimation. Therefore, we can only derive similar theoretical guarantees with ZOEG with improved dependence on the dimension. When $\tau_{x}=d_{x}$, and $\tau_{y}=d_{y}$, ZOBCEG estimates full gradients and recovers the ZOCEG algorithm. However, our subsequent numerical results show that ZOBCEG can typically achieve comparable oracle complexities to ZOCEG even when $\tau_{x}<d_{x}$ and $\tau_{y}<d_{y}$.
Algorithm 3 Zeroth-Order Block Coordinate Extra-Gradient (ZOBCEG)
  Inputs: Number of iterations $K$, step size $\eta$, smoothing radii $\{r_{k}\}$, and initial point $\mathbf{z}_{0}$.
  for $k\xleftarrow{}0$ to $K-1$ do
     Randomly sample $\mathcal{I}_{k}^{x}\subseteq\{1,2,\cdots,d_{x}\}$, $\mathcal{I}_{k}^{x}\subseteq\{1,2,\cdots,d_{x}\}$, with $|\mathcal{I}_{k}^{x}|=\tau_{x}$, $|\mathcal{I}_{k}^{y}|=\tau_{y}$, and calculate $\mathbf{z}_{k}^{+}$ by:
$\mathbf{x}^{+}_{k}=\mathcal{P}_{\mathcal{X}}\left[\mathbf{x}_{k}-\eta\tilde{\mathbf{g}}_{k}^{x}\right],$
$\mathbf{y}^{+}_{k}=\mathcal{P}_{\mathcal{Y}}\left[\mathbf{y}_{k}+\eta\tilde{\mathbf{g}}_{k}^{y}\right].$
Based on $\mathbf{z}_{k}^{+}$, randomly sample $\mathcal{J}_{k}^{x}\subseteq\{1,2,\cdots,d_{x}\}$, $\mathcal{J}_{k}^{x}\subseteq\{1,2,\cdots,d_{x}\}$, with $|\mathcal{J}_{k}^{x}|=\tau_{x}$, $|\mathcal{J}_{k}^{y}|=\tau_{y}$, and update $\mathbf{z}_{k}$ by:
$\mathbf{x}_{k+1}=\mathcal{P}_{\mathcal{X}}\left[\mathbf{x}_{k}-\eta\tilde{\mathbf{g}}_{k}^{x,+}\right],$
$\mathbf{y}_{k+1}=\mathcal{P}_{\mathcal{Y}}\left[\mathbf{y}_{k}+\eta\tilde{\mathbf{g}}_{k}^{y,+}\right].$
  end for
  Outputs: $\hat{\mathbf{z}}_{K}=\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{z}^{+}_{k}.$
5   Numerical simulations
In this section, we conduct numerical experiments to validate the performance of the proposed algorithms on the demand-side load tracking problem in power systems in both convex and nonconvex test cases, as in [17]. An aggregator tries to coordinate the loads of multiple users to curtail the total power load $p_{c}(\mathbf{x})$ and track the target load level while minimizing the total cost $\phi_{0}(\mathbf{x})$. The problem is formulated as follows:
$\min_{\mathbf{x}\in\mathcal{X}}\ \phi_{0}(\mathbf{x}),$
$s.t.\ \ p_{c}(\mathbf{x})\leq D,$
where $p_{c}(\mathbf{x})$ is determined by the dynamics of the distribution system and seen as a black box.
5.1   Convex case
In the convex case, we apply a simplified linear model of $p_{c}(x):\;p_{c}(\mathbf{x})=\sum_{i=1}^{100}(1+\gamma_{i})(u_{i}-x_{i})$, where $\mathbf{x}$ denotes the load adjustment vector. We consider $\phi_{0}(\mathbf{x})=\sum_{i=1}^{100}h_{i}(x_{i})$ and $h_{i}(x_{i})$ is the local cost function of user $i$, which is modeled as $h_{i}(x_{i})=a_{i}x_{i}^{2}+b_{i}x_{i}$. The feasible set is defined as $\mathcal{X}=\left\{\mathbf{x}\in\mathbb{R}^{100}\mid 0\leq x_{i}\leq u_{i},i=1,...,100\right\}$.
In addition to testing our proposed algorithms, we also compare them with two other algorithms, stochastic zeroth-order constraint extrapolation (SZO-ConEX) [30] and zeroth-order gradient descent ascent (ZOGDA) [20], applicable to solving problem (1). Both of them employ the two-point gradient estimator. We run the tests with constant step size (CS) and diminishing step size (DS), respectively, where the DS is set as $\eta_{k}:=\eta_{0}/\sqrt{k+1}$. Moreover, we evaluate the performance of ZOBCEG under different block sizes (BS) $\tau_{x}$. Note that setting $BS=100$ recovers the ZOCEG algorithm, which adopts full gradient estimation.
Regarding the experimental setup, the coefficients $a_{i}$ and $b_{i}$ in the local cost functions are sampled from the uniform distribution $a_{i}\sim U(0.5,1.5)$ and $b_{i}\sim U(0,5)$, respectively. The initial load levels $u_{i},\forall i$ are randomly selected from $U(0\text{kW},50\text{kW})$. As for the constraint, each coefficient $\gamma_{i}$ is independently sampled from $U(0.03,0.15)$, and the desired load level is set to $D=(p_{c}(\mathbf{0})-1500)\text{kW}$. In all tests, we use the smoothing radius of $r_{k}=\min\{5/(k+1)^{1.1},10^{-3}\}$, and run each test 20 times with different initial points. We test all algorithms with multiple choices of step sizes, and the simulation results with the best step sizes are shown in Fig 1 and Fig 2, where the relative error is computed by $(\phi_{0}(\mathbf{x}_{k})-\phi_{0}^{*})/\phi_{0}^{*}$. Here, $\phi_{0}^{*}$ is the optimal function value derived by an oracle solver with explicit formulation of problem (1). The dark lines represent average values, and the shaded areas represent standard deviations.
Figure 1: Performance comparison in the convex case: (a) average relative error under CS, (b) average constraint violation under CS, (c) average relative error under DS, (d) average constraint violation under DS.
As shown in Fig 1, ZOEG can converge with proper step sizes and achieve satisfactory performance in both objective optimality and constraint violation. All tested algorithms benefit from diminishing step sizes, which consistently improve the convergence behaviors. Among them, SZO-ConEX enjoys faster convergence in constraint violation but performs worse in relative error. In contrast, ZOEG and ZOGDA achieve similar and competitive performance with faster convergence in relative error. The comparable performance of ZOEG and ZOGDA aligns with the iteration complexity proportional to $\epsilon^{-2}$ of stochastic GDA [28], which has a close relation to ZOGDA.
Figure 2: Performance of ZOBCEG in the convex case: (a) average relative error, (b) average constraint violation.
Figure 2 illustrates the performance of ZOBCEG across varying BSs. Overall, ZOBCEG achieves faster convergence and smaller relative errors compared to the above algorithms. Convergence rates improve with larger BSs. Notably, ZOCEG (i.e., ZOBCEG with BS=100) converges within a few dozen iterations, aligning with our theoretical analysis and confirming improved convergence rates and computational efficiency. While ZOBCEG’s theoretical complexity bounds are unfavorable, it demonstrates strong practical performance. Furthermore, Table 2 reports the average number of function evaluations required to reach specific target errors. Although ZOCEG exhibits a fast convergence rate, its requirement of $2d$ function evaluations per iteration incurs a higher single-step computational cost. In contrast, ZOBCEG with an appropriate BS offers a favorable trade-off between evaluations per iteration and total computational complexity, which can potentially reduce overall cost.
Table 2: Average number of function evaluations required to achieve specific errors under different block sizes.
BS Relative error Constraint violation
5% 1% 0.10% 5 1 0.1
1 2460.6 4247.1 5664.9 210.6 359.7 1309.2
5 905.8 1479.1 1786.4 183.4 466.2 1488.9
100 581.4 1458.6 2723.4 2152.2 2876.4 4324.8
5.2   Nonconvex case
For the nonconvex case, we consider the demand-side energy optimization at a feeder of the 141-bus distribution network. The objective function includes the penalty for load and voltage deviation, along with the local load management costs: $\phi_{0}(\mathbf{x})=20\cdot(p_{c}(\mathbf{x})-D)^{2}+20\cdot\sum_{j=1}^{141}\left(\max\{v_{j}(\mathbf{x})-\bar{v},0\}^{2}+\max\{\underline{v}-v_{j}(\mathbf{x}),0\}^{2}\right)+\sum_{i}^{168}h_{i}(x_{i})$. The total power load $p_{c}(\mathbf{x})$ and voltage magnitude $v_{j}$ at bus $j$ capture the nonlinear and nonconvex nature of AC power flows and serve as black boxes. $\mathbf{x}$ includes active and reactive power loads across all buses. and the function $h_{i}$ and the set $\mathcal{X}$ are generated as in the convex setting. We evaluate the performance of ZOEG under CS and DS, and ZOBCEG with various BS under CS, and adopt the technique in [17] to ensure the feasibility when perturbing $\mathbf{x}$.
In this case, we set parameters $\bar{v}=1.04$ p.u., $\underline{v}=0.96$ p.u., and $D=0.15$ p.u. $=1500kW$. For ZOEG, we set $\eta=3\times 10^{-6}$, $\delta=0.005$ for CS and $\eta_{k}=3\times 10^{-4}/\sqrt{k+1000}$, $\delta_{k}=\min\{50/(k+1),0.1\}$ for DS. The smoothing radius is $r_{k}=\min\{0.01/(k+4000)^{1.1},10^{-5}\}$. For ZOBCEG, we consider CS with different BS, and the smoothing radius is $r_{k}=\min\{0.1/k^{1.2},2\times 10^{-4}\}$.
We run each case 50 times with different initial points, and the results of ZOEG and ZOBCEG are shown in Figure 3. The proposed algorithms demonstrate excellent convergence performance even in nonconvex settings, highlighting their effectiveness in addressing more complex optimization problems. Moreover, despite having similar theoretical complexity bounds in the convex case, ZOBCEG exhibits a faster empirical convergence rate than ZOEG, and the convergence performance of ZOBCEG improves as the BS increases.
Figure 3: Performance of algorithms in the nonconvex case: (a) average objective value of ZOEG, (b) average constraint violation of ZOEG, (c) average objective value of ZOBCEG, (d) average constraint violation of ZOBCEG.
6   Conclusions
This work addresses a general constrained optimization problem characterized by black-box objective and constraint functions. To tackle the challenge of lacking analytical formulations, we reformulate it as a min-max problem and introduce two novel zeroth-order algorithms: the zeroth-order extra-gradient (ZOEG), leveraging a 2-point gradient estimator, and the zeroth-order coordinate extra-gradient algorithm (ZOCEG), based on a $2d$-point coordinate gradient estimator. ZOCEG is further extended to a block-coordinate variant, the zeroth-order block coordinate extra-gradient (ZOBCEG), which updates randomly selected coordinate blocks. Theoretical analysis establishes finite-sample convergence guarantees for both ZOEG and ZOCEG in convex settings regarding both optimality gap and constraint violation. Notably, ZOEG improves dependence on the problem dimension, while ZOCEG achieves the best-known convergence rate. Numerical experiments validate the efficiency of these algorithms in both convex and nonconvex scenarios.
There are several promising directions for future research: 1) Extending the algorithms to handle stochastic settings where black-box observations are corrupted by noise; 2) Improving the theoretical complexity analysis of ZOBCEG, which currently lags behind its empirical performance; 3) Establishing the theoretical guarantees for the proposed methods in nonconvex settings.
Appendices
Appendix A Proof of Lemma 2.1
By definition, we have
$\ \left|G_{h,i}(\mathbf{z};r)-\nabla_{i}h(\mathbf{z})\right|$
$=$ $\ \left|\frac{h(\mathbf{z}+r\mathbf{e}_{i})-h(\mathbf{z})}{r}-\nabla_{i}h(\mathbf{z})\right|$
$=$ $\ \left|\int_{0}^{1}\nabla_{i}h(\mathbf{z}+s\cdot r\mathbf{e}_{i})-\nabla_{i}h(\mathbf{z}_{k})ds\right|$
[…]
where the second inequality follows from the smoothness.
A.1   Proof of Lemma 3.1
Firstly, by introducing $\left\langle\eta F^{r_{k}}(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle$, we have:
$\ \left\langle\eta F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle$
$=$ $\ \mathbb{E}_{\mathbf{w}_{k}}\left[\left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]+\left\langle\eta F(\mathbf{z}_{k}^{+})-\eta F^{r_{k}}(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle$
$\leq$ $\ \mathbb{E}_{\mathbf{w}_{k}}\left[\left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]+\eta\|F(\mathbf{z}_{k}^{+})-F^{r_{k}}(\mathbf{z}_{k}^{+})\|\|\mathbf{z}_{k}^{+}-\mathbf{z}\|$
$\leq$ $\ \mathbb{E}_{\mathbf{w}_{k}}\left[\left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]+\eta Lr_{k}\cdot\tilde{D},$
where the first equality follows from (6) and the last inequality follows from (7). Taking expectation over $\mathbf{v}_{k}$ gives:
$\ \mathbb{E}\left[\left\langle\eta F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\mid\mathcal{F}_{k}\right]$ (16)
$\leq$ $\ \mathbb{E}\left[\left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\mid\mathcal{F}_{k}\right]+\eta L\tilde{D}r_{k}.$
Then, we focus on the bound for the first term:
$\ \left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle$
$=$ $\ \left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\right\rangle+\left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k+1}-\mathbf{z}\right\rangle$
$\leq$ $\ \left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\right\rangle-\left\langle\mathbf{z}_{k+1}-\mathbf{z}_{k},\mathbf{z}_{k+1}-\mathbf{z}\right\rangle$
$=$ $\ \left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\right\rangle-\frac{1}{2}\|\mathbf{z}_{k+1}-\mathbf{z}_{k}\|^{2}+H_{k}(\mathbf{z})-H_{k+1}(\mathbf{z}),$ (17)
where the inequality follow from the projection property:
$\left\langle\eta\mathbf{g}_{k}^{+}+\mathbf{z}_{k+1}-\mathbf{z}_{k},\mathbf{z}_{k+1}-\mathbf{z}\right\rangle\leq 0.$
Similarly, by introducing the auxiliary term $\langle\eta\mathbf{g}_{k},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\rangle$, we can obtain the upper bound for $\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\rangle-\frac{1}{2}\|\mathbf{z}_{k+1}-\mathbf{z}_{k}\|^{2}$ as follows:
$\ \left\langle\eta\mathbf{g}_{k}^{+},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\right\rangle-\frac{1}{2}\|\mathbf{z}_{k+1}-\mathbf{z}_{k}\|^{2}$
$=$ $\ \eta\left\langle\mathbf{g}_{k}^{+}-\mathbf{g}_{k},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\right\rangle+\left\langle\eta\mathbf{g}_{k},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\right\rangle-\frac{1}{2}\|\mathbf{z}_{k+1}-\mathbf{z}_{k}\|^{2}$
$\leq$ $\ \eta\left\langle\mathbf{g}_{k}^{+}-\mathbf{g}_{k},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\right\rangle-\left\langle\mathbf{z}_{k}^{+}-\mathbf{z}_{k},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\right\rangle-\frac{1}{2}\|\mathbf{z}_{k+1}-\mathbf{z}_{k}\|^{2}$
$=$ $\ \eta\left\langle\mathbf{g}_{k}^{+}-\mathbf{g}_{k},\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\right\rangle-\frac{1}{2}\|\mathbf{z}_{k}^{+}-\mathbf{z}_{k}\|^{2}-\frac{1}{2}\|\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\|^{2}$
$\leq$ $\ \eta\|\mathbf{g}_{k}-\mathbf{g}_{k}^{+}\|\|\mathbf{z}_{k}^{+}-\mathbf{z}_{k+1}\|$
$\leq$ $\ \eta^{2}\|\mathbf{g}_{k}-\mathbf{g}_{k}^{+}\|^{2},$ (18)
where the first inequality uses the projection property:
$\left\langle\eta\mathbf{g}_{k}+\mathbf{z}_{k}^{+}-\mathbf{z}_{k},\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\leq 0,$
and the final bound follows from Cauchy-Schwarz inequality and the contraction property of the projection operator.
Since
$\mathbf{g}_{k}-\mathbf{g}_{k}^{+}=\mathbf{g}_{k}-F(\mathbf{z}_{k})+F(\mathbf{z}_{k})-F(\mathbf{z}_{k}^{+})+F(\mathbf{z}_{k}^{+})-\mathbf{g}_{k}^{+},$
we can bound the conditional expectation of $\|\mathbf{g}_{k}-\mathbf{g}_{k}^{+}\|^{2}$:
$\ \mathbb{E}\left[\|\mathbf{g}_{k}-\mathbf{g}_{k}^{+}\|^{2}\mid\mathcal{F}_{k}\right]$
$\leq$ $\ 3\cdot\mathbb{E}\left[\|\mathbf{g}_{k}-F(\mathbf{z}_{k})\|^{2}+\|\mathbf{g}_{k}^{+}-F(\mathbf{z}_{k}^{+})\|^{2}\mid\mathcal{F}_{k}\right]+3\cdot\mathbb{E}\left[\|F(\mathbf{z}_{k})-F(\mathbf{z}_{k}^{+})\|^{2}\mid\mathcal{F}_{k}\right]$
$\leq$ $\ 3\cdot\mathbb{E}\left[\|\mathbf{g}_{k}-F(\mathbf{z}_{k})\|^{2}\mid\mathcal{F}_{k}\right]+3\cdot\mathbb{E}\left[\|\mathbf{g}_{k}^{+}-F(\mathbf{z}_{k}^{+})\|^{2}\mid\mathcal{F}_{k}\right]+12G^{2}.$
Note that combining the bound for variance of gradient estimator (7) with (8) yields the following bound:
$\mathbb{E}\left[\|\mathbf{g}_{k}-F(\mathbf{z}_{k})\|^{2}\mid\mathcal{F}_{k}\right]$
$\leq$ $2\cdot\mathbb{E}\left[\|\mathbf{g}_{k}-F^{r_{k}}(\mathbf{z}_{k})\|^{2}+\|F^{r_{k}}(\mathbf{z}_{k})-F(\mathbf{z}_{k})\|^{2}\mid\mathcal{F}_{k}\right]$
$\leq$ $\frac{18d^{2}G^{2}}{d+2}+\frac{3d^{2}L^{2}r_{k}^{2}}{2}+2L^{2}r_{k}^{2},$
and the bound on $\mathbb{E}\left[\|\mathbf{g}_{k}^{+}-F(\mathbf{z}_{k}^{+})\|^{2}\mid\mathcal{F}_{k}\right]$ can be derived similarly. Then, $\mathbb{E}\left[\|\mathbf{g}_{k}-\mathbf{g}_{k}^{+}\|^{2}\mid\mathcal{F}_{k}\right]$ is bounded as follows:
$\mathbb{E}\left[\|\mathbf{g}_{k}-\mathbf{g}_{k}^{+}\|^{2}\mid\mathcal{F}_{k}\right]\leq\hat{\Delta}_{k},$ (19)
where $\hat{\Delta}_{k}=\frac{108d^{2}G^{2}}{d+2}+9d^{2}L^{2}r_{k}^{2}+12L^{2}r_{k}^{2}+12G^{2}$. The proof is completed by combining the bounds established in (19), (A.1), (A.1), and (LABEL:eq:_pf_bound_cond_first_order).
Appendix B Proof of Lemma 4.1
For notation simplicity, we denote $H_{k}^{x}(\mathbf{x})=\frac{1}{2}\|\mathbf{x}_{k}-\mathbf{x}\|^{2}$, and $H_{k}^{y}(\mathbf{y})=\frac{1}{2}\|\mathbf{y}_{k}-\mathbf{y}\|^{2}$. First, we deal with $\left\langle\eta\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$ by introducing $\eta\left\langle\hat{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$:
$\ \left\langle\eta\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$
$=$ $\ \eta\left\langle\hat{\mathbf{g}}_{k}^{x,+}+\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+})-\hat{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$
$\leq$ $\ \eta\left\langle\hat{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle+\frac{1}{2}\eta Lr_{k}\sqrt{d_{x}}\tilde{D},$ (20)
where the last inequality follows from the bounded bias of coordinate gradient estimator: $\|\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+})-\hat{\mathbf{g}}_{k}^{x,+}\|\leq\frac{\sqrt{d_{x}}}{2}Lr_{k}$, which directly follows from (5).
We then focus on the bound on the first term $\eta\left\langle\hat{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$ as follows:
$\ \eta\left\langle\hat{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$
$=$ $\ \eta\left\langle\hat{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle+\eta\left\langle\hat{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k+1}-\mathbf{x}\right\rangle$
$\leq$ $\ \eta\left\langle\hat{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle-\left\langle\mathbf{x}_{k+1}-\mathbf{x}_{k},\mathbf{x}_{k+1}-\mathbf{x}\right\rangle$
$=$ $\ \eta\left\langle\hat{\mathbf{g}}_{k}^{x,+}-\hat{\mathbf{g}}_{k}^{x},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle+\eta\left\langle\hat{\mathbf{g}}_{k}^{x},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle-\left\langle\mathbf{x}_{k+1}-\mathbf{x}_{k},\mathbf{x}_{k+1}-\mathbf{x}\right\rangle$
$\leq$ $\ \eta\left\langle\hat{\mathbf{g}}_{k}^{x,+}-\hat{\mathbf{g}}_{k}^{x},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle-\left\langle\mathbf{x}_{k}^{+}-\mathbf{x}_{k},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle-\left\langle\mathbf{x}_{k+1}-\mathbf{x}_{k},\mathbf{x}_{k+1}-\mathbf{x}\right\rangle,$ (21)
where the two inequalities follow from the properties of projection:
$\left\langle\eta\hat{\mathbf{g}}_{k}^{x,+}+\mathbf{x}_{k+1}-\mathbf{x}_{k},\mathbf{x}_{k+1}-\mathbf{x}\right\rangle\leq 0,$
$\left\langle\eta\hat{\mathbf{g}}_{k}^{x}+\mathbf{x}_{k}^{+}-\mathbf{x}_{k},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\leq 0.$
By performing a simple algebraic transformation, we have:
$\ \left\langle\mathbf{x}_{k}-\mathbf{x}_{k}^{+},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle+\left\langle\mathbf{x}_{k}-\mathbf{x}_{k+1},\mathbf{x}_{k+1}-\mathbf{x}\right\rangle$ (22)
$=$ $\ H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})-\frac{1}{2}\|\mathbf{x}_{k+1}-\mathbf{x}_{k}^{+}\|^{2}-\frac{1}{2}\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k}\|^{2}.$
Besides, $\left\langle\hat{\mathbf{g}}_{k}^{x,+}-\hat{\mathbf{g}}_{k}^{x},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle$ can be bounded as follows:
$\ \left\langle\hat{\mathbf{g}}_{k}^{x,+}-\hat{\mathbf{g}}_{k}^{x},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle$
$=$ $\ \left\langle\hat{\mathbf{g}}_{k}^{x,+}-\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+})+\nabla_{\mathbf{x}}f(\mathbf{z}_{k})-\hat{\mathbf{g}}_{k}^{x},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle+\left\langle\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+})-\nabla_{\mathbf{x}}f(\mathbf{z}_{k}),\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle$
$\leq$ $\ Lr_{k}\sqrt{d_{x}}\tilde{D}+L\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k}\|\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\|+L\|\mathbf{y}_{k}^{+}-\mathbf{y}_{k}\|\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\|,$ (23)
where the last step follows from Lemma 2.1 and the smoothness of $f$. By the condition $\eta L\leq\frac{1}{2}$, substituting (LABEL:eq:_full_grad_alge), (B), and (B) into (B) gives:
$\ \left\langle\eta\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$
$\leq$ $\ \frac{3}{2}\eta Lr_{k}\sqrt{d_{x}}\tilde{D}+\eta L\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k}\|\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\|+H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})-\frac{1}{2}\|\mathbf{x}_{k+1}-\mathbf{x}_{k}^{+}\|^{2}$
$\ -\frac{1}{2}\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k}\|^{2}+\eta L\|\mathbf{y}_{k}^{+}-\mathbf{y}_{k}\|\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\|$
$\leq$ $\ H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})+\frac{3}{2}\eta Lr_{k}\sqrt{d_{x}}\tilde{D}-\frac{1}{4}\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k}\|^{2}+\frac{1}{4}\|\mathbf{y}_{k}^{+}-\mathbf{y}_{k}\|^{2}.$ (24)
where in the last step we used $a\cdot b\leq\frac{1}{2}(a^{2}+b^{2}),\forall a,b\in\mathbb{R}$. Similarly, we have:
$-\left\langle\eta\nabla_{\mathbf{y}}f(\mathbf{z}_{k}^{+}),\mathbf{y}_{k}^{+}-\mathbf{y}\right\rangle\leq H_{k}^{y}(\mathbf{y})-H_{k+1}^{y}(\mathbf{y})+\frac{3}{2}\eta Lr_{k}\sqrt{d_{y}}\tilde{D}-\frac{1}{4}\|\mathbf{y}_{k}^{+}-\mathbf{y}_{k}\|^{2}+\frac{1}{4}\|\mathbf{x}_{k}^{+}-\mathbf{x}_{k}\|^{2}.$ (25)
Adding (B) and (25) gives:
$\left\langle\eta F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\leq H_{k}(\mathbf{z})-H_{k+1}(\mathbf{z})+\frac{3}{2}\eta Lr_{k}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}.$
Appendix C Proof of Theorem 4.3
Similarly, We start by defining a filtration $\mathcal{F}_{k}^{b}:=\sigma\left(\mathbf{z}_{0},\mathcal{I}_{0}^{x},\mathcal{I}_{0}^{y},\mathcal{J}_{0}^{x},\mathcal{J}_{0}^{y},\mathbf{z}_{1},...,\mathbf{z}_{k}\right)$. Given $\mathbf{z}_{k}$, $\mathbf{z}_{k}^{+}$, $\langle\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\allowbreak\mathbf{x}_{k}^{+}-\mathbf{x}\rangle$ is bounded as follows:
$\ \left\langle\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$
$\leq$ $\ \left\langle\mathbb{E}_{\mathcal{J}_{k}^{x}}\left[\frac{d_{x}}{\tau_{x}}\tilde{\mathbf{g}}_{k}^{x,+}\right],\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle+\left\|\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+})-\mathbb{E}_{\mathcal{J}_{k}^{x}}\left[\frac{d_{x}}{\tau_{x}}\tilde{\mathbf{g}}_{k}^{x,+}\right]\right\|\cdot\|\mathbf{x}_{k}^{+}-\mathbf{x}\|$
$\leq$ $\ \left\langle\mathbb{E}_{\mathcal{J}_{k}^{x}}\left[\frac{d_{x}}{\tau_{x}}\tilde{\mathbf{g}}_{k}^{x,+}\right],\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle+\frac{Lr_{k}\sqrt{d_{x}}\tilde{D}}{2}.$
Then, taking expectation over $\mathcal{I}_{k}^{x},\mathcal{I}_{k}^{y}$ gives:
$\ \mathbb{E}\left[\left\langle\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\mid\mathcal{F}_{k}^{b}\right]$ (26)
$\leq$ $\ \mathbb{E}\left[\left\langle\frac{d_{x}}{\tau_{x}}\tilde{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\mid\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}\sqrt{d_{x}}\tilde{D}}{2}.$
Specifically, for $\left\langle\tilde{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$, we have:
$\ \left\langle\tilde{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle$
$\leq$ $\ \left\langle\tilde{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle-\frac{1}{\eta}\left\langle\mathbf{x}_{k+1}-\mathbf{x}_{k},\mathbf{x}_{k+1}-\mathbf{x}\right\rangle$
$\leq$ $\ \left\langle\tilde{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta}$
$\leq$ $\ \eta\|\tilde{\mathbf{g}}_{k}^{x,+}\|\cdot\|\tilde{\mathbf{g}}_{k}^{x,+}-\tilde{\mathbf{g}}_{k}^{x}\|+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta}$
$\leq$ $\ \eta\left(2G^{2}+2GLr_{k}\sqrt{\tau_{x}}+\frac{L^{2}r_{k}^{2}\tau_{x}}{2}\right)+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta},$ (27)
where the first inequality follows from the projection property, the second one follows from algebraic transformation, and the last inequality follows from $\|\tilde{\mathbf{g}}_{k}^{x,+}\|\leq G+\frac{Lr_{k}\sqrt{\tau_{x}}}{2}$. Substituting (C) into (LABEL:eq:_block_grad_bound_ZOCEG_term) gives:
$\ \mathbb{E}\left[\left\langle\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\mid\mathcal{F}_{k}^{b}\right]$ (28)
$\leq$ $\ \frac{d_{x}}{\eta\tau_{x}}\mathbb{E}\left[H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})\mid\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}\sqrt{d_{x}}\tilde{D}}{2}+\frac{\eta d_{x}}{\tau_{x}}\left(2G^{2}+2GLr_{k}\sqrt{\tau_{x}}+\frac{L^{2}r_{k}^{2}\tau_{x}}{2}\right).$
Similarly, for the dual variable, there holds:
$\ \mathbb{E}\left[-\left\langle\nabla_{\mathbf{y}}f(\mathbf{z}_{k}^{+}),\mathbf{y}_{k}^{+}-\mathbf{y}\right\rangle\mid\mathcal{F}_{k}^{b}\right]$ (29)
$\leq$ $\ \frac{d_{y}}{\eta\tau_{y}}\mathbb{E}\left[H_{k}^{y}(\mathbf{y})-H_{k+1}^{y}(\mathbf{y})\mid\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}\sqrt{d_{y}}\tilde{D}}{2}+\frac{\eta d_{y}}{\tau_{y}}\left(2G^{2}+2GLr_{k}\sqrt{\tau_{y}}+\frac{L^{2}r_{k}^{2}\tau_{y}}{2}\right).$
Combining (LABEL:eq:_block_grad_bound_x_term) and (LABEL:eq:_block_grad_bound_y_term) gives:
$\ \mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\mid\mathcal{F}_{k}^{b}\right]$
$\leq$ $\ \frac{Lr_{k}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2}+\frac{d_{x}}{\eta\tau_{x}}\mathbb{E}\left[H_{k}^{x}(\mathbf{z})-H_{k+1}^{x}(\mathbf{z})\mid\mathcal{F}_{k}^{b}\right]+\frac{d_{y}}{\eta\tau_{y}}\mathbb{E}\left[H_{k}^{y}(\mathbf{z})-H_{k+1}^{y}(\mathbf{z})\mid\mathcal{F}_{k}^{b}\right]$
$\ +2\eta G^{2}\left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)+2\eta GLr_{k}(\frac{d_{x}}{\sqrt{\tau_{x}}}+\frac{d_{y}}{\sqrt{\tau_{y}}})+\frac{d_{x}+d_{y}}{2}\eta L^{2}r_{k}^{2}.$
Taking the total expectation of and calculate the telescoping sum for $k=0,1,\cdots,K-1$, we can bound $\sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]$ as follows:
$\ \sum_{k=0}^{K-1}\mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right]$
$\leq$ $\ \frac{LM_{1}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2}+\frac{d_{x}}{\eta\tau_{x}}\mathbb{E}\left[H_{0}^{x}(\mathbf{z})-H_{K}^{x}(\mathbf{z})\mid\mathcal{F}_{k}^{b}\right]+\frac{d_{y}}{\eta\tau_{y}}\mathbb{E}\left[H_{0}^{y}(\mathbf{z})-H_{K}^{y}(\mathbf{z})\mid\mathcal{F}_{k}^{b}\right]$
$\ +2\eta KG^{2}\left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)+2\eta GLM_{1}(\frac{d_{x}}{\sqrt{\tau_{x}}}+\frac{d_{y}}{\sqrt{\tau_{y}}})+\frac{d_{x}+d_{y}}{2}\eta L^{2}M_{2}$
$\leq$ $\ \frac{LM_{1}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2}+\frac{d_{x}\tilde{D}^{2}}{2\eta\tau_{x}}+\frac{d_{y}\tilde{D}^{2}}{2\eta\tau_{y}}+2\eta KG^{2}\left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)+2\eta GLM_{1}(\frac{d_{x}}{\sqrt{\tau_{x}}}+\frac{d_{y}}{\sqrt{\tau_{y}}})$
$\ +\frac{d_{x}+d_{y}}{2}\eta L^{2}M_{2},$ (30)
where the last inequality follows from the boundedness of $\mathcal{Z}$. Substituting (C) into Lemma 3.2 finishes the convergence analysis of ZOBCEG.